OpenStudy (anonymous):

what are the 1st and 2nd derivatives of X^2/((x^2)+1)

OpenStudy (anonymous):

Use Quotient Rule....

OpenStudy (anonymous):

\[\frac{ d }{ dx }(\frac{ x^2 }{ x^2+1 })\]

OpenStudy (anonymous):

Yeah, I don't get that either

OpenStudy (anonymous):

\[=\frac{ (x^2+1) \frac{ d }{ dx }x^2 - x^2\frac{ d }{ dx }(x^2+1) }{ (x^2+1)^2 }\]